OpenStudy (anonymous):

If A is the set of all numbers a^2+4ab+b^2, where a and b are both integers, prove that A is closed under multiplication. I've tried a few different things, but the algebra always gets out of hand. Any help would be appreciated!

OpenStudy (perl):

so we want to show the set of numbers of the form p is closed under multiplication where p= a^2 +4ab + b^2 and a,b are integers. Now to show closure we need two elements. call them p1 and p2 . show that p1 * p2 is of the form m^2 + 4m*n + n^2 for some integers m,n

OpenStudy (anonymous):

Got there, and multiplied it out, but I couldn't see a way forward.

OpenStudy (anonymous):

Also, wouldn't we need to fix m and n instead? The choice of m and n is arbitrary, but we should need to fix them to prove closure.

OpenStudy (perl):

now you need two integers for each number. so let p1 = r^2 + 4rs + s^2 and p2 = u^2 + 4u*v+ v^2

OpenStudy (anonymous):

*Sorry, didn't see you were still typing

OpenStudy (perl):

then p1 * p2 = (r^2 + 4rs + s^2)*(u^2 + 4u*v + v^2)

OpenStudy (perl):

= r^2*u^2+4*r^2*u*v+r^2*v^2+4*r*s*u^2+16*r*s*u*v+4*r*s*v^2+s^2*u^2+4*s^2*u*v+s^2*v^2

OpenStudy (anonymous):

Looks like what I got

OpenStudy (perl):

can we simplify this

OpenStudy (perl):

we can rewrite this sum r^2*u^2+4*r^2*u*v+r^2*v^2+4*r*s*u^2+16*r*s*u*v+4*r*s*v^2+s^2*u^2+4*s^2*u*v+s^2*v^2 = r^2u^2 + r^2v^2 + s^2u^2 + s^2v^2 + 4( r^2uv + rsu^2 + rsv^2 + s^2uv + 4rsuv)

OpenStudy (anonymous):

That's what I got also (using different variables, so it's taking me a couple minutes to check)

OpenStudy (anonymous):

WolframAlpha says that bit inside the parentheses isn't factorable, so that's where I got stuck.

OpenStudy (anonymous):

Unless some of those terms belong with the square terms (the a^2 and b^2 terms)

OpenStudy (perl):

yes this looks pretty ugly , you can say that the product of two integers is an integer, and keep using this result

OpenStudy (perl):

but i dont see the exact way to get it in the form m^2 + 4mn + n^2

OpenStudy (anonymous):

Me either. I generated the first few numbers, and it seems to check out.

OpenStudy (perl):

there is an identity you can use a^2+b^2+4ab=(a+2b)^2-3b^2

OpenStudy (perl):

that might be easier to deal with

OpenStudy (anonymous):

I tried these guys: \[(a+b)^2 +2ab\]\[2(a+b)-(a^2+b^2)\]

OpenStudy (perl):

so using the identity, we have p1 * p2 = (r^2 + 4rs + s^2)*(u^2 + 4u*v + v^2) = [ (r + 2s)^2 - 3s^2] [ (u + 2v)^2 - 3v^2]

OpenStudy (anonymous):

Looks good

OpenStudy (perl):

so using the identity, we have p1 * p2 = (r^2 + 4rs + s^2)*(u^2 + 4u*v + v^2) = [ (r + 2s)^2 - 3s^2] [ (u + 2v)^2 - 3v^2] = (r+2s)^2*(u+2v)^2 - 3s^2*(u+2v)^2 - 3v^2*(r+2s)^2 + 9s^2v^2

OpenStudy (perl):

ok i found a new identity, this one will work :)

OpenStudy (perl):

if f(x,y) = x^2 + 4xy + y^2 Then f(a,b)* f(c,d)=f(ac-bd,ad+4bd+bc)

OpenStudy (perl):

do you see how it works now?

OpenStudy (perl):

f(r,s) * f(u,v) = f( ru - sv , rv + 4sv + su ) so those will be our m and n

OpenStudy (perl):

(r^2 + 4rs + s^2)*(u^2 + 4u*v + v^2) = (ru-sv)^2 + 4 (ru - sv) *(rv +4sv + su) + (rv + 4sv + su)^2

OpenStudy (anonymous):

In your initial setup for the function, how did you get your x to be ac-bd?

OpenStudy (perl):

now let m = (ru-sv) , n = (rv + 4sv + su)

OpenStudy (perl):

sorry, i didnt understand

OpenStudy (anonymous):

f(a,b)* f(c,d)=f(ac-bd,ad+4bd+bc) Where did you get the ac-bd part?

OpenStudy (perl):

but it works

OpenStudy (anonymous):

Identity for which part? I see that it's supposed to be x, but ? = ac-bd

OpenStudy (perl):

break it down into parts remember the rule is f(x,y) = x^2 + 4xy + y^2 so... f(a,b) = a^2+4ab+b^2 f(c,d) = c^2+4cd+d^2 f(ac-bd,ad+4bd+bc) = (ac-bd)^2 + 4(ac-bd)(ad+4bd+bc) + (ad+4bd+bc)^2

OpenStudy (anonymous):

Ah, alright, back with you now.

OpenStudy (perl):

so we are basically using the identity (a^2 +4ab + b^2)(c^2+4cd+d^2)= (ac-bd)^2 + 4(ac-bd)(ad+4bd+bc) + (ad+4bd+bc)^2

OpenStudy (perl):

the lefthand side can be rewritten using the shorthand f(a,b) * f(c,d) the right side can be written using the shorthand f( ac-bd, ad+4bd + bc)

OpenStudy (anonymous):

I see the identity now :) . Awesome. How did you find that?

OpenStudy (perl):

if the f(a,b) stuff confuses you , you can ignore it. the main thing is the identity (a^2 +4ab + b^2)(c^2+4cd+d^2)= (ac-bd)^2 + 4(ac-bd)(ad+4bd+bc) + (ad+4bd+bc)^2 we can write the product of two numbers of the set A as again a member of the set of A.

OpenStudy (anonymous):

No, it doesn't. Just thought that you derived it somehow and I didn't see where it comes from. I can see that it works, but I'm wondering how it was derived in the first place.

OpenStudy (perl):

oh sorry about that :)

OpenStudy (anonymous):

Nah, it's awesome. I have been making 0 headway with this problem. How did you find that identity?

OpenStudy (perl):

yeah i dont see an easy way to derive this I found it here http://math.stackexchange.com/questions/644861/if-both-integers-x-and-y-can-be-represented-as-a2-b2-4ab-prove-that see the comment by heropup

OpenStudy (anonymous):

Thanks :)

OpenStudy (perl):

scroll down to the second answer

OpenStudy (perl):

you can use a more abstract approach, using norms what class are you in

OpenStudy (anonymous):

I'm a middle school teacher (I was a math major), but a friend was asking me about a few problems for some Stanford math camp. They're allowed to use whatever they want as long as they say where they got help. Thanks a lot for the help!

OpenStudy (perl):

ahh, ok :)

OpenStudy (perl):

I tutor math for a living, this was a fun problem :)

OpenStudy (perl):

oh it looks like there is another way to do this, or perhaps it is the same

OpenStudy (anonymous):

I'm pretty sure that ac-bd thing came from the idea of a normed vector space. I've seen it before, but I'm not sure where. I think it's used a lot in programming...

OpenStudy (perl):

we have two identities: a^2+b^2+4ab=(a+2b)^2-3b^2 And : (x^2-3y^2)(u^2-3v^2)=(xu+3yv)^2-3(xv+yu)^2

OpenStudy (perl):

so to do this requires two substitutions, keeping track of everything is kind of tricky though

OpenStudy (anonymous):

Yeah, I'm going to throw this at her and make her do it :P

OpenStudy (perl):

r^2 + 4rs + s^2 = ( r + 2s ) ^2 - 3s^2 u^2 + 4uv + v^2 = (u + 2v)^2 - 3v^2

OpenStudy (perl):

:D

OpenStudy (perl):

i was doing this before, but i didnt finish it

OpenStudy (anonymous):

I'll let her play with it and I may repost if she doesn't get anywhere.

OpenStudy (anonymous):

These problems are pretty fun. There's one where I'm wondering if they gave enough information.

OpenStudy (perl):

sure :)

OpenStudy (perl):

but of course to get this identity simply from thin air probably isnt going to happen, so you have to use those other more well known tricks/identities

OpenStudy (anonymous):

Yeah, that looks like what they did on here with the Brahmagupta-Fibonacci Identity

OpenStudy (perl):

are you from USA?

OpenStudy (anonymous):

yes, VA You?

OpenStudy (perl):

yes from new york

OpenStudy (anonymous):

Nice. Get a lot of snow these past few weeks?

OpenStudy (perl):

yes , hehe